OpenStudy (anonymous):

the shorter leg of a right triangle if 6 centimeters less than the other leg.Find the length of the two legs if the hypotenuse is 30 centimeters

OpenStudy (anonymous):

x^2 + (x+6)^2 = 30^2 solve

OpenStudy (anonymous):

You know that a^2+b^2=c^2 when a and b are the two sides and c is the hypotenuse and that leads you to sourwings' equation :)